# Searching for strong lensing by late-type galaxies in UNIONS

### v1

@@ 4.3 Reparameterising the mass models
[…]
For every potential deflector, we create a regular grid of parameters:
  • $z_{\rm s}\in\{\mathrm{e}^{x}\mid x\text{ uniformly sampled between}\ln(z_{\rm d})\text{ and}\ln(3.0)\}$
  • $\Upsilon\in\{4,5,6,7,8,9\}$
  • $f_{\ast}\in\{0.4,0.48,0.56,0.64,0.72,0.8\}$
For every combination of parameters in the above grid, $z_{\rm s},\Upsilon,$ and $f_{\ast}$; we find the corresponding combination of $z_{\rm s}$, $\alpha_{1}$, and $M_{200}$; such that $\kappa(z_{d},z_{s},\boldsymbol{L},\Upsilon,f_{\ast})=\kappa(z_{\rm d},z_{\rm s},\boldsymbol{L},\alpha_{1},M_{200})$. Then, we compute $\log(M)$ and $\theta_{\mathrm{E}}$. For these searches we use the bisection method as is implemented in scipy (Virtanen et al. 2020). For each deflector, we save all the compatible values of the seven parameters: $z_{\rm s},\Upsilon,f_{\ast},\alpha_{1},\log(M_{200}),\log(M)$, and $\theta_{\mathrm{E}}$. Following that, we estimate the values in-between the grid sampling using the scikit-learn (Pedregosa et al. 2011) implementation of the ‘Random Sample Consensus’ algorithm with a linear regression as base regressor (Fischler & Bolles 1981). We use $\Upsilon$, $f_{\ast}$, and $\theta_{\mathrm{E}}$ to predict $z_{\rm s}$, $\alpha_{1}$, $M_{200}$, and $\log(M)$. For this regression, we normalise the data set such that each parameter has a mean of 0 and a standard deviation of 1, and generate polynomial features using a polynomial of degree 3. We take two-thirds of the sampled models for training and leave the remaining one-third for testing. We keep the interpolated values only if the accuracy on the test set is larger than 98%. This has to be done independently for each deflector, since the relationship between the seven parameters depends on the redshift of the deflector and its luminosity. The output of this procedure is a well populated grid of reasonable mass models, such that for every potential deflector and every combination of physical parameters $\Upsilon$, $f_{\ast}$, $\log(M)$, and $\theta_{\mathrm{E}}$, we have the corresponding $z_{\rm s}$ and the Lenstronomy parameters $\alpha_{1}$, and $M_{200}$.
@@ 4.4 Generating a mock lens from a potential deflector
[…]
@@
We randomly select between the tangential and radial caustics and place the central pixel of the source inside a polygon with the same shape as the selected caustic, but with a 15% larger perimeter. We favour the selection of the tangential over the radial caustic with a probability of 75% versus 25%. Before simulating the lensing of the source by the convergence of the deflector, we interpolate the light of the source from the HST image and rotate it by a random angle. The lensed image is generated with a pixel size exactly 6 times smaller than that of the deflector. We convolve the high resolution lensed source image with the corresponding PSF model of the deflector. Then, we downsample the resulting image to the CFIS pixel size of $0.1857″$ and add Poissonian noise. The final mock lens is the sum of the deflector with the simulated image of the lensed source.
At this stage all the simulated lenses are physically motivated images mocking CFIS observations of gravitational lensing. However, the lensing features can be too faint or too blended with the deflector light, up to the point that it becomes impossible for human experts to discern the lensing features. This is often the case when simulating small Einstein radius systems and it also affects known lenses when they are seen at worse resolution. For example, the SLACS lenses (Bolton et al. 2006), which were discovered spectroscopically and thus are rather compact, are not identified as lenses by human experts when seen on the Dark Energy Survey (DES) imaging (Rojas et al. 2023). Consequently, even if there are arcs or multiple images after simulating the lensing, that by itself is not enough to guarantee that the lensing features are visible and identifiable as such in the final image. This is especially problematic when simulating lensing by edge-on late-type galaxies because their light profiles tend to be extended and to have substructure; thus, increasing the chances of blending between the deflector light and the lensed source. We address this problem by implementing a ‘contrast check’ to detect if the lensing features are visible in the final image. If a mock fails the contrast check, we change the position of the source relative to the caustic lines and generate a new mock lens. If the new mock also fails the contrast check, we reject it and proceed to the next potential deflector. The details of the contrast check are presented in Appendix A. Finally, we show a sample of 56 mock lenses with different Einstein radius in Appendix B.
@@ 5 Finding lenses with CMU DeepLens
Figure 5: Left: histogram in logarithmic scale of the distribution of network scores for the full parent sample, only 17 5141751417\,51417 514 of the sources had a score larger than 0.9. Right: precision and recall versus network score for the test data set. The dashed line marks the classification threshold applied on the parent sample, 0.9. We balance the precision, recall, and number of sources selected to settle on a classification threshold.
We use the CNN CMU DeepLens (Lanusse et al. 2018) for the binary classification task. This is a residual neural network (ResNet, see He et al. 2016a, b) explicitly designed for lens finding. It is the best performing entry in “The strong gravitational lens finding challenge” (Metcalf et al. 2019) and a modified version has been successfully used in multiple lens finding efforts (Huang et al. 2020, 2021; Storfer et al. 2024; Shu et al. 2022). We reimplement the network and train it from scratch using the PyTorch framework (Paszke et al. 2019).
@@ 5.1 Training
[…]
We split the data set into three parts: a training set, with 80% of the sources; and validation and test sets, each with 10% of the sources. The subsets are kept balanced, with an equal number of lenses and non-lenses. During training, each image goes through the following data processing steps:
@@
    random rotation up to 180°180°180  ° using a bilinear interpolation;
@@
    cropping down to 55 pixels with a new central pixel randomly selected up to 11 pixels away from the previous centre, effectively shifting the image up to 22″2  ″;
@@
    vertical flip with a 50% chance;
@@
    normalisation to mean 0 and standard deviation 1.
During inference, the images are still cropped to 55 pixels and normalised, but the central pixel remains fixed.
We train the model by optimising a cross-entropy loss function using a stochastic gradient descent (SGD) optimiser with an initial learning rate of $10^{-3}$ and a momentum of $0.9$. The weights of the network are updated using the loss calculated on the training data, while monitoring the loss on the validation set at the end of every training epoch. If the loss on the validation set does not improve for 20 consecutive epochs, then the learning rate is reduced to half its current value, down to a minimum learning rate of $10^{-8}$, at which the learning rate is kept constant. The training ends after 50 consecutive epochs without improvement of the validation loss, training for a total for 350 epochs, with the best validation loss being at the 300th epoch. The batch size during training is 512 and the images are shuffled at every epoch. Overall, the training takes 17 hours on an NVIDIA GeForce GTX 1080 Ti graphics card.
@@ 5.2 Metrics and inference
Figure 6: Edge-on lens candidates identified by the trained CNN, or discovered serendipitously either by prototypes of the classifier or while preparing the mock lenses. They correspond to five grade A, 20 grade B, and 59 grade C; of which only two are already known. The layout follows the same scheme from Fig. 2.
Figure 7: Non-edge-on lens candidates identified by the trained CNN, or discovered serendipitously either by prototypes of the classifier or while preparing the mock lenses. The layout follows the same scheme from Fig. 2. The 38 candidates correspond to 22 grade A, 12 grade B, and four grade C. Of which, 12 are reported in SLED.
Figure 8: Follow-up observations for J1121+4351 and J1559+3147. Left: UNIONS $r$-band and $u$-band imaging of the lens candidates. The red dashed line represents the slit position. Right: reduced 1D spectra from long-slit spectroscopy taken with the ALFOSC instrument in the NOT. The grey-shaded area marks atmospheric absorption bands.
@@
We monitor the evolution of the loss function on both the training and validation data sets. This is shown in the left panel of Fig. 4. The trends of these two sets exhibit no significant divergence, suggesting that overfitting on the training set is not a concern. In the right panel of Fig. 4, we present the receiver operating characteristic (ROC) curve for the training, validation, and test sets. The performance of the algorithm is remarkably consistent across the three data sets, with a marginally better performance on the training set, although, this difference is negligible. Furthermore, the performance of the trained network on the test and validation data sets is, at first glance, very good, with an area under the ROC curve of 0.9992, very close to the optimal 1.0. However, this performance is underestimating the false positive rate when applied to the parent sample. Because of the extremely low prevalence of edge-on lenses, estimated to be 1 in 30 000 in Sect. 3, the classifier must have a false positive rate of the order of 1/30 0001300001/30\,0001 / 30 000 for the sample of CNN-selected lens candidates to have a purity of 50%. Failing this, the network will suggest hundreds of false positives per actual edge-on lens candidate.
@@
We apply the trained network to the full parent sample, yielding a network score for each of the 6 978 97769789776\,978\,9776 978 977 sources. The distribution of these scores is presented in the left panel of Fig. 5. We use the distribution of scores, together with the precision and recall curves evaluated on the test set, to choose a reasonable classification threshold for the parent sample. This is shown in the right panel of Fig. 5. We select a classification threshold of 0.9, which balances a tolerable number of false positives with the potential number of lenses to be discovered. In total, the network suggests 17 5141751417\,51417 514 potential candidates.
[…]
@@ 5.3 Visual inspection of CNN candidates
Before visually inspecting the sample selected by the CNN, we removed the candidates discovered during the characterisation of the parent sample, that is, the sources presented in Fig. 2. We then applied the same three-step visual inspection methodology introduced in Sect. 3. The first two steps of the visual inspection were performed by J.A.A.B., B.C., and F.C., while the final group inspection was conducted by J.A.A.B., B.C., F.C., C.L., R.G., and K.R. After the first two steps, 107 sources were selected for the group inspection. To that sample, we added lens candidates discovered while prototyping our implementation of CMU Deeplens, as well as serendipitous candidates discovered while preparing the simulations used to train the network. In total, 144 sources were inspected by the six experts in the final 1-by-1 group inspection.
@@ 6 NOT spectroscopy
[…]
@@
The selected lens candidates are: J1559+3147, a grade A edge-on lens candidate previously reported in Cañameras et al. (2020) and Savary et al. (2022); and J1121+4351, a newly discovered grade B edge-on candidate. Both candidates have existing SDSS spectra, which provide the redshift of the deflectors. However, although the potential lensing features of J1121+4351 are within the SDSS fibre, the SDSS model residuals lack sufficient signal-to-noise to constrain the redshift of the source. By contrast, the brightest lensing features of J1559+3147 are outside the SDSS fibre aperture. Overall, they are well-suited candidates for confirmation with single-epoch long-slit spectroscopy focused on the lensing features.
@@
We reduced the observations and extracted calibrated 1D spectra following the methodology of Lemon et al. (2023) and Dux et al. (2024). In brief, the 2D spectra were bias-subtracted and cosmic rays masked using a threshold in the absolute Laplacian. A spatially-resolved background was then estimated and removed using the median value around the trace, and an error map was generated by combining background and Poissonian noise. A wavelength solution was obtained by fitting HeNe arc exposures, and refined by fitting against sky lines. The 1D spectra were extracted by dividing the trace into discrete wavelength bins and fitting a Moffat profile to each source in the trace (3 for J1559+3147 and 2 for J1121+4351) at every wavelength bin. The parameters of the Moffat profile were then interpolated using a polynomial fit, enabling accurate recovery of the 1D spectra even if the trace shifts spatially. Finally, the 1D spectra were derived by evaluating the Moffat profiles at every wavelength. Following this step, we applied flux calibration using the spectrum of the standard star Feige 34. The resulting 1D spectra are presented in Fig. 8.
@@
Regarding the measurement of redshifts, we find for the deflector galaxy in J1121+4351 a series of absorption lines: Ca H and K, Fe G, Mg, and Na, which allows us to confirm the SDSS measurement of $z_{\rm d}=0.097$. For the potential background source, we detect the emission lines [O II] and [O III], indicating star formation and a redshift of $z_{\rm s}=0.246$. This higher redshift is consistent with the lensing hypothesis, suggesting that the source is a lensed arc. However, the lack of detection of a counter image, combined with the ground-based resolution of CFHT, prevents us from making a definitive confirmation. Higher resolution imaging is required. Promisingly, the Euclid Wide Survey will provide high-resolution observations for most of the lens candidates discovered in this work. This is crucial for edge-on lenses since they have a fair chance of being in a naked-cusp configuration, leading to the absence of a radial counter image, and thus requiring high-resolution imaging to better identify tangential lensed features. In the case of J1559+3147, we observe that the blue spectra share identical features, confirming that they are in fact two images of a single source. Furthermore, we identify [O ii] in emission and K in absorption, which we use to obtain a redshift measurement: $z_{\rm s}=0.691$. Also, we identify Ca H and K, Mg, and Na in absorption in the spectrum of the deflector, confirming the SDSS redshift: $z_{\rm d}=0.149$. Remarkably, neither deflector shows emission lines.

### v2

@@ 4.3 Reparameterising the mass models
However, in order to avoid ‘unphysical’ or unreal-looking lenses, we impose several constraints on the mass models. More precisely, we force the stellar mass-to-light ratio to be in the range $4.0<\Upsilon<9.0$, the fraction of stellar to total mass to be $0.4<f_{\ast}<0.8$ (baryonic fraction), and the total mass to be $10^{10}\,<\,M/M_{\sun}\,<\,10^{12}$. All three quantities are calculated in the area enclosed by the tangential critical lines. In order to do this efficiently we reparameterise the total convergence from $\kappa\left(z_{\rm d},z_{\rm s},\mathbf{L},\alpha_{1},M_{200}\right)$ to $\kappa\left(z_{\rm d},z_{\rm s},\mathbf{L},\Upsilon,f_{\ast}\right)$.
For every potential deflector, we created a regular grid of parameters:
  • $z_{\rm s}\in\{\mathrm{e}^{x}\mid x\text{ uniformly sampled between}\ln(z_{\rm d})\text{ and}\ln(3.0)\}$,
  • $\Upsilon\in\{4,5,6,7,8,9\}$, and
  • $f_{\ast}\in\{0.4,0.48,0.56,0.64,0.72,0.8\}$.
For every combination of parameters in the above grid, $z_{\rm s},\Upsilon,$ and $f_{\ast}$; we find the corresponding combination of $z_{\rm s}$, $\alpha_{1}$, and $M_{200}$; such that $\kappa(z_{d},z_{s},\boldsymbol{L},\Upsilon,f_{\ast})=\kappa(z_{\rm d},z_{\rm s},\boldsymbol{L},\alpha_{1},M_{200})$. Then, we compute $\log(M)$ and $\theta_{\mathrm{E}}$. For these searches we use the bisection method as is implemented in scipy (Virtanen et al. 2020). For each deflector, we save all the compatible values of the seven parameters: $z_{\rm s},\Upsilon,f_{\ast},\alpha_{1},\log(M_{200}),\log(M)$, and $\theta_{\mathrm{E}}$. Following that, we estimate the values in-between the grid sampling using the scikit-learn (Pedregosa et al. 2011) implementation of the ‘Random Sample Consensus’ algorithm with a linear regression as base regressor (Fischler and Bolles 1981). We use $\Upsilon$, $f_{\ast}$, and $\theta_{\mathrm{E}}$ to predict $z_{\rm s}$, $\alpha_{1}$, $M_{200}$, and $\log(M)$. For this regression, we normalise the dataset such that each parameter has a mean of 0 and a standard deviation of 1, and generate polynomial features using a polynomial of degree 3. We take two-thirds of the sampled models for training and leave the remaining one-third for testing. We keep the interpolated values only if the accuracy on the test set is larger than 98%. This has to be done independently for each deflector, since the relationship between the seven parameters depends on the redshift of the deflector and its luminosity. The output of this procedure is a well populated grid of reasonable mass models, such that for every potential deflector and every combination of physical parameters $\Upsilon$, $f_{\ast}$, $\log(M)$, and $\theta_{\mathrm{E}}$, we have the corresponding $z_{\rm s}$ and the Lenstronomy parameters $\alpha_{1}$, and $M_{200}$.
@@ 4.4 Generating a mock lens from a potential deflector
There is evidence that ML classifiers benefit from a ‘flat’ Einstein radius distribution when training on simulated lenses (Cañameras et al. 2024). Consequently, we produce the same number of mocks for every target Einstein radius, thus ensuring a flat $\theta_{\mathrm{E}}$ distribution between 1.2 and 2.6, with a precision of 0.1.
@@
We randomly select between the tangential and radial caustics and place the central pixel of the source inside a polygon with the same shape as the selected caustic, but with a 15% larger perimeter. We favour the selection of the tangential over the radial caustic with a probability of 75% versus 25%. Before simulating the lensing of the source by the convergence of the deflector, we interpolate the light of the source from the HST image and rotate it by a random angle. The lensed image is generated with a pixel size exactly 6 times smaller than that of the deflector. We convolve the high resolution lensed source image with the corresponding PSF model of the deflector. Then, we downsample the resulting image to the CFIS pixel size of 0.1857 and add Poissonian noise. The final mock lens is the sum of the deflector with the simulated image of the lensed source.
At this stage all the simulated lenses are physically motivated images mocking CFIS observations of gravitational lensing. However, the lensing features can be too faint or too blended with the deflector light, up to the point that it becomes impossible for human experts to discern the lensing features. This is often the case when simulating small Einstein radius systems and it also affects known lenses when they are seen at worse resolution. For example, the SLACS lenses (Bolton et al. 2006), which were discovered spectroscopically and thus are rather compact, are not identified as lenses by human experts when seen on the Dark Energy Survey (DES) imaging (Rojas et al. 2023). Consequently, even if there are arcs or multiple images after simulating the lensing, that by itself is not enough to guarantee that the lensing features are visible and identifiable as such in the final image. This is especially problematic when simulating lensing by edge-on late-type galaxies because their light profiles tend to be extended and to have substructure, thus increasing the chances of blending between the deflector light and the lensed source. We address this problem by implementing a ‘contrast check’ to detect if the lensing features are visible in the final image. If a mock fails the contrast check, we change the position of the source relative to the caustic lines and generate a new mock lens. If the new mock also fails the contrast check, we reject it and proceed to the next potential deflector. The details of the contrast check are presented in Appendix C. Finally, we show a sample of 56 mock lenses with different Einstein radius in Appendix D, and a characterisation of the mock population in Appendix E.
@@ 5 Finding lenses with CMU DeepLens
Figure 5: Left: Histogram in logarithmic scale of the distribution of network scores for the full parent sample. Only 17 51417\,514 of the sources had a score greater than 0.9. Right: Precision and recall versus network score for the test dataset. The dashed line marks the classification threshold applied on the parent sample, 0.9. We balanced the precision, recall, and number of sources selected to settle on a classification threshold.
Figure 6: Follow-up observations for J1121+4351 and J1559+3147. Left: UNIONS $r$-band and $u$-band imaging of the lens candidates. The red dashed line represents the slit position. Right: Reduced 1D spectra from long-slit spectroscopy taken with the ALFOSC instrument in the NOT. The grey-shaded area marks atmospheric absorption bands. Spectral lines are highlighted with vertical dashed lines.
We used the CNN CMU DeepLens (Lanusse et al. 2018) for the binary classification task. This is a residual neural network (ResNet, see He et al. 2016a, b) explicitly designed for lens finding. It is the best performing entry in ‘The strong gravitational lens finding challenge’ (Metcalf et al. 2019) and a modified version has been successfully used in multiple lens finding efforts (Huang et al. 2020, 2021; Storfer et al. 2024; Shu et al. 2022). We reimplemented the network and trained it from scratch using the PyTorch framework (Paszke et al. 2019).
@@ 5.1 Training
The dataset consists of 150 000150\,000 cutouts of 66× $6\,\times\,66$ pixels with a pixel size of 0.1857. The dataset is balanced, with 75 00075\,000 simulated lenses following the prescriptions of Sect. 4, and an equal number of non-lenses. The simulated lenses are distributed uniformly across a range of Einstein radii between 1.2 and 2.6, with exactly $5000$ mocks per 0.1 bin. We present in Appendix D a mosaic with some of the mock lenses in this sample. For the non-lenses, we reuse the stamps inspected in Sect. 3 and keep a flat distribution in magnitude by randomly selecting 18 75018\,750 sources per magnitude bin. By doing so, we guarantee the absence of potential lens candidates in the negative training set. We present a characterisation of the image properties of the training data in Appendix E.
We split the dataset into three parts: a training set, with 80% of the sources, and validation and test sets, each with 10% of the sources. The subsets are kept balanced, with an equal number of lenses and non-lenses. During training, each image goes through the following data processing steps:
@@
    Random rotation up to 180° using a bilinear interpolation;
@@
    Cropping down to 55 pixels with a new central pixel randomly selected up to 11 pixels away from the previous centre, effectively shifting the image up to 2;
@@
    Vertical flip with a 50% chance; and
@@
    Normalisation to a mean of zero and a standard deviation of one.
During inference, the images were still cropped to 55 pixels and normalised, but the central pixel remained fixed.
We trained the model by optimising a cross-entropy loss function using a stochastic gradient descent optimiser with an initial learning rate of $10^{-3}$ and a momentum of $0.9$. The weights of the network are updated using the loss calculated on the training data, while monitoring the loss on the validation set at the end of every training epoch. If the loss on the validation set does not improve for 20 consecutive epochs, then the learning rate is reduced to half its current value, down to a minimum learning rate of $10^{-8}$, at which the learning rate is kept constant. The training ends after 50 consecutive epochs without improvement of the validation loss, training for a total for 350 epochs, with the best validation loss being at the 300th epoch. The batch size during training is 512 and the images are shuffled at every epoch. Overall, the training takes 17 hours on a NVIDIA GeForce GTX 1080 Ti graphics card.
@@ 5.2 Metrics and inference
We monitored the evolution of the loss function on both the training and validation datasets. This is shown in the left panel of Fig. 4. The trends of these two sets exhibit no significant divergence, suggesting that overfitting on the training set is not a concern. In the right panel of Fig. 4, we present the receiver operating characteristic (ROC) curve for the training, validation, and test sets. The performance of the algorithm is remarkably consistent across the three datasets, with a marginally better performance on the training set, although, this difference is negligible. Furthermore, the performance of the trained network on the test and validation datasets is, at first glance, very good, with an area under the ROC curve of 0.9992, very close to the optimal 1.0. However, these metrics underestimate the FPR when applied to the parent sample. Because of the extremely low prevalence of edge-on lenses, estimated to be 1 in 30 000 in Sect. 3, the classifier must have a FPR of the order of 1/30 0001/30\,000 for the sample of CNN-selected lens candidates to have a purity of 50%. Failing this, the network will suggest hundreds of false positives per actual edge-on lens candidate.
@@
We applied the trained network to the full parent sample, yielding a network score for each of the 6 978 9776\,978\,977 sources. The distribution of these scores is presented in the left panel of Fig. 5. We use the distribution of scores, together with the precision and recall curves evaluated on the test set, to choose a reasonable classification threshold for the parent sample. This is shown in the right panel of Fig. 5. We select a classification threshold of 0.9, which balances a tolerable number of false positives with the potential number of lenses to be discovered. In total, the network suggests 17 51417\,514 potential candidates.
Lastly, we used Eq. 3 to estimate the number of good lens candidates selected by the network. This requires estimates of the model’s recall and FPR, as well as the general prevalence of edge-on lenses in the parent sample. The recall of the model is estimated using the mock lenses in the test set, corresponding to $P(\text{C}\,|\,\text{L})=0.95$. Similarly, the FPR is estimated from the non-lenses in the test set. However, this subset contains an equal number of elements in each of the magnitude bins defined in Sect. 3, and thus we calculated the FPR per bin. The weighted average of these values, using the total number of elements per bin in the parent sample as weights, gives the overall FPR, corresponding to $P(\text{C}\,|\,\text{Non-L})=2.6\times 10^{-3}$. Finally, using the edge-on lens prevalence from Sect. 3, $P(L)\approx 1/30\,000$, and Eq. 3, we predict that $P(\text{L}\,|\,\text{C})\approx 1.2\%$ of the candidates selected by the network will be good candidates. This corresponds to about 208 good edge-on lens candidates in the 17 51417\,514 machine-selected sample. Since both the FPR and the prevalence of edge-on lensing are estimated directly from the data, any deviation from this prediction is most likely due to the estimation of the recall, which is based on mock simulations.
@@ 5.3 Visual inspection of CNN candidates
Before visually inspecting the sample selected by the CNN, we removed the candidates discovered during the characterisation of the parent sample, that is, the sources presented in Fig. 2. We then applied the same three-step visual inspection methodology introduced in Sect. 3. The first two steps of the visual inspection were performed by J.A.A.B., B.C., and F.C., while the final group inspection was conducted by J.A.A.B., B.C., F.C., C.L., R.G., and K.R. After the first two steps, 107 sources were selected for the group inspection. To that sample, we added lens candidates discovered while prototyping our implementation of CMU Deeplens, as well as serendipitous candidates discovered while preparing the simulations used to train the network. In total, 144 sources were inspected by the six experts in the final one-by-one group inspection.
@@ 6 NOT spectroscopy
We obtained long-slit spectroscopic follow-up for two lens candidates using the ALFOSC instrument at NOT over two consecutive nights: 14 and 15 of April, 2023. For each lens candidate, we acquired two $1200\,\text{s}$ observations using grism 4 with a 1.0-wide slit, providing a spectral resolution of $3.3\,\text{\AA }\,\text{pixel}^{-1}$ and a wavelength coverage spanning $3200\,\text{\AA }$ to $9600\,\text{\AA }$. The average seeing during the observations was 0.9.
@@
The selected lens candidates are: J1559+3147, a grade A edge-on lens candidate previously reported in Cañameras et al. (2020) and Savary et al. (2022); and J1121+4351, a newly discovered grade B edge-on candidate. Both candidates have existing SDSS spectra, which provide the redshift of the deflectors. However, although the potential lensing features of J1121+4351 are within the SDSS fibre, the SDSS model residuals lack sufficient signal-to-noise ratio (S/N) to constrain the redshift of the source. By contrast, the brightest lensing features of J1559+3147 are outside the SDSS fibre aperture. Overall, they are well-suited candidates for confirmation with single-epoch long-slit spectroscopy focused on the lensing features.
@@
We reduced the observations and extracted calibrated 1D spectra following the methodology of Lemon et al. (2023) and Dux et al. (2024). In brief, the 2D spectra were bias-subtracted and cosmic rays masked using a threshold in the absolute Laplacian. A spatially resolved background was then estimated and removed using the median value around the trace, and an error map was generated by combining background and Poissonian noise. A wavelength solution was obtained by fitting HeNe arc exposures, and refined by fitting against sky lines. The 1D spectra were extracted by dividing the trace into discrete wavelength bins and fitting a Moffat profile to each source in the trace (3 for J1559+3147 and 2 for J1121+4351) at every wavelength bin. The parameters of the Moffat profile were then interpolated using a polynomial fit, enabling accurate recovery of the 1D spectra even if the trace shifts spatially. Finally, the 1D spectra were derived by evaluating the Moffat profiles at every wavelength. Following this step, we applied flux calibration using the spectrum of the standard star Feige 34. The resulting 1D spectra are presented in Fig. 6.
@@
Regarding the measurement of redshifts, we find for the deflector galaxy in J1121+4351 a series of absorption lines: Ca H and K, Fe G, Mg, and Na, which allows us to confirm the SDSS measurement of $z_{\rm d}=0.097$. For the potential background source, we detect the emission lines [O II] and [O III], indicating star formation and a redshift of $z_{\rm s}=0.246$. This higher redshift is consistent with the lensing hypothesis, suggesting that the source is a lensed arc. However, the lack of detection of a counter image, combined with the ground-based resolution of CFHT, prevents us from making a definitive confirmation. Higher resolution imaging is required. Promisingly, the Euclid Wide Survey will provide high-resolution observations for most of the lens candidates discovered in this work. This is crucial for edge-on lenses since they have a fair chance of being in a naked-cusp configuration, leading to the absence of a radial counter image and thus requiring high-resolution imaging to better identify tangential lensed features. In the case of J1559+3147, we observe that the blue spectra share identical features, confirming that they are in fact two images of a single source. Furthermore, we identify [O ii] in emission and K in absorption, which we use to obtain a redshift measurement: $z_{\rm s}=0.691$. Also, we identify Ca H and K, Mg, and Na in absorption in the spectrum of the deflector, confirming the SDSS redshift: $z_{\rm d}=0.149$. Remarkably, neither deflector shows emission lines.